OpenStudy (anonymous):

find the solutions to the system y=x^2 + 5x +6 y=-4x+12

hartnn (hartnn):

solving them simultaneously , eliminating y x^2 +5x+6 = -4x+12 this is a quadratic in x, can you solve this ?

OpenStudy (anonymous):

when i try i come up with the wrong awnser

hartnn (hartnn):

what values of x you got ?

OpenStudy (anonymous):

2,20 and -3,0

hartnn (hartnn):

x^2 +5x+6 = -4x+12 x^2+9x -6 = 0 and 2,-3 are certainly not the solutions for x can you try to solve x^2+9x -6 = 0 again ? are you sure question is correct ?as there are no integer solutions for x here

OpenStudy (anonymous):

like the awnser choices are 2,-20 and -3,0 2,20 and -3, -18 -2,-20 and -3,-18 no soultions so i guse i should pick D?

hartnn (hartnn):

if the question is correct, then yes, D it is.....

hartnn (hartnn):

but i'd say even D is not correct if D was 'no integral solution' , then that would be correct

OpenStudy (anonymous):

no the four in y=4x+12 is positive not negative

hartnn (hartnn):

ohh! so, the quadratic becomes x^2 +x -6 = 0 this is easy to solve, can you ?

OpenStudy (anonymous):

i can't say that i can

hartnn (hartnn):

2 numbers with sum +1 and product = -6 are ... ?

OpenStudy (anonymous):

-2 + 3?

hartnn (hartnn):

yes so we split x as +3x-2x x^2+3x-2x -6 x (x+3) - 2 (x+3) can you further factor out x+3 ?

OpenStudy (anonymous):

no because it is an odd number?

hartnn (hartnn):

you can factor out that term \((x+3)(x-2\))

OpenStudy (anonymous):

how?

hartnn (hartnn):

because that term was common in both the terms! :)

OpenStudy (anonymous):

so that would be the awnser ?

hartnn (hartnn):

so, x = 2,-3

hartnn (hartnn):

and since y = 4x+12 find y when x= 2 find y when x= -3

OpenStudy (anonymous):

thanks bro